with a tail wind, a helicopter traveled 300 miles in 1 hour 20 minutes. The return trip against the same wind took 20 minutes longer. Find the plane's air speed and the speed of the helicopter.
Answer by venugopalramana(3286)

SEE THE FOLLOWING EXAMPLE AND TRY
-----------------------------------------
With a head wind, a plane traveled 1000 miles in 4 hours. With the same wind as a tail wind, the return trip took 3 hours 20 minutes. Find the plane's air speed( P SAY ) and the wind speed ( W SAY)
LET DISTANCE TRAVELLED = 1000 MILES
RELATIVE SPEED WHILE GOING HEAD WIND = P-W MPH
TIME NEEDED = T = D/R.S = 1000/(P-W) = 4
P-W=250.................1
R.S. WHILE GOING TAIL WIND = P+W
TIME NEEDED = 1000/(P+W)=3 HRS 20 MTS = 3+20/60 = 10/3 HRS
P+W =300.......................2
EQN.1+EQN.2
2P=550
P=275 MPH
W = 300-275 = 25 MPH